### v1

# Thermalization rates and quantum Ruelle-Pollicott resonances: insights from operator hydrodynamics

Abstract
In thermalizing many-body quantum systems without conservation laws such as ergodic Floquet and random unitary circuits, local expectation values are predicted to decay to their equilibrium values exponentially quickly. In this work we derive a relationship between said exponential decay rate $\overline{g}$ and the operator spreading properties of a local unitary evolution. A hydrodynamical picture for operator spreading allows us to argue that, for random unitary circuits, $\overline{g}$ is encoded by the leading eigenvalue of a dynamical map obtained by enriching unitary dynamics with dissipation, in the limit of weak dissipation. We argue that the size of the eigenvalue does not depend on the details of this weak dissipation (given mild assumptions on properties of the ergodic dynamics), so long as it only suppresses large operators significantly. Our calculations are based on analytical results for random unitary circuits, but we argue that similar results hold for ergodic Floquet systems. These conjectures are in accordance with existing results which numerically obtain quantum many-body analogues of classical Ruelle-Pollicott resonances [T. Prosen J. Phys. A: Math. Gen. 35 L737 (2002), T. Mori, arXiv:2311.10304] by studying unitary evolutions subject to weak dissipation.
Introduction.— In a thermalizing many-body system, local expectation values decay towards their equilibrium ensemble averages, irrespective of most of the details of the initial state. In a system without any global symmetries, such as an ergodic Floquet and random unitary circuit (RUC), this decay is expected to be exponential Prosen (2002, 2004); von Keyserlingk et al. (2018); Nahum et al. (2018); Mori et al. (2018); Lucas (2019); Schiulaz et al. (2019); Mori (2024). This is, via linear response theory, tied to the expected exponential decay of autocorrelation functions of local operators
$|\langle O(t)O(0)\rangle-\langle O(t)\rangle\langle O(0)\rangle|\sim e^{-\bar{g}t}.$ (1)
Here $\langle\cdot\rangle=q^{-L}\mathrm{Tr}(\cdot)$ is the infinite temperature ensemble expectation value, suitable for ergodic quantum systems without symmetry. $L$ denotes system size, and $q$ is the on-site Hilbert space dimension. Eq. (1) is expected to hold until time scales proportional to $L$.
The decay rate $\bar{g}$ captures how quickly an ergodic system thermalizes, and thus locally loses information about its initial state. [^1] Note that it is distinct from the butterfly velocity $v_{B}$: while $v_{B}$ describes how quickly operators grow apart, $\bar{g}$ describes how how slowly an operator returns, and these two rates can be very different. We discuss this point more in the Discussion and the supplemental material sup . Our work here concerns three connected questions: 1) How can $\bar{g}$, which describes loss of information, be understood from the underlying unitary dynamics? 2) $e^{-\bar{g}}$ is not an eigenvalue of a unitary operator such as the Floquet unitary. Can this quantity be encoded as an eigenvalue of a related operator? 3) Is $\bar{g}$ sensitive to the choice of operator $O$ or the initial state of the system?
To answer the first question, we leverage the hydrodynamical description for operator growth introduced in Refs. Nahum et al., 2018; von Keyserlingk et al., 2018. In short, the information loss is closely related to the fact that, under time evolution, local operators rapidly grow in spatial support, hiding the detailed local information in the initial state into ever more delicate many-body correlators. The hydrodynamical description for operator growth can be packaged into a stochastic matrix with a leading eigenvalue 1 due to unitarity. Roughly speaking, its second leading eigenvalue can be identified with $e^{-\bar{g}}$ (there are some subtleties, which we will address later). From the hydrodynamical description, as we will show, it is also straightforward to answer question 3 in the negative, in systems without symmetry.
Question 2 was explored in classical dynamical systems, where $\bar{g}$ is related to Ruelle-Pollicott (RP) resonances Ruelle (1986); Pollicott (1985), which are given by poles of matrix elements of the analytically continued resolvent $R(z)=(z-U)^{-1}$, located at eigenvalues of $U$. Here, $U$ is the unitary Frobenius-Perron operator generating the discrete (Floquet) dynamics on the phase space densityGaspard (2005). These poles are difficult to compute exactly outside of simple models Hasegawa and Saphir (1992). A useful approach for computing these resonances numerically is to coarse-grain the phase space, so that $U$ becomes non-unitary Weber et al. (2000); Klus et al. (2015). Then $U$ can have eigenvalues $\{e^{-i\epsilon_{j}}\}$ within the unit circle. It has been shown in certain cases that the largest eigenvalue is “frozen” (independent of the coarse graining, in a certain limit), coinciding with the leading RP resonance which sets the decay rate of autocorrelations Nonnenmacher (2003); Faure and Roy (2006)
Figure 1: $(a)$ Ref. Mori, 2024 suggested obtaining $\bar{g}$ for unitary evolution (pink rectangles) by adding a small amount of dissipation (black squares), and taking the system size $L$ to infinity before taking the dissipation rate to zero. $(b)$ In a hydrodynamical description of operator evolution, where $\overline{\rho_{R}}(x,t)$ describes the density of operators with right endpoint at site $x$ at time $t$, the unitary evolution is described by a stochastic matrix $T$ that tends to grow operators while the dissipation is described by a diagonal matrix $P(\gamma)$ that suppresses operators with large weight/size.
More recently, various authors have studied quantum analogues of RP resonances, by applying quantum versions of “coarse-graining of phase space” Prosen (2002, 2004); Mori (2024); Znidaric (2024). Ref. Prosen, 2002, 2004 considered adding a projection $\mathbb{P}_{\ell}$ to an ergodic Floquet system, so that the Heisenberg dynamics of an operator $O$ is given by $\mathbb{P}_{\ell}\mathbb{U}_{F}$. Here $\mathbb{U}_{F}[\cdot]=U_{F}^{\dagger}[\cdot]U_{F}$ generates the Heisenberg dynamics under the Floquet unitary $U_{F}$. $\mathbb{P}_{\ell}$ is a dissipative superoperator that projects out operators with support greater than $\ell$. The resulting superoperator $\mathbb{P}_{\ell}\mathbb{U}_{F}$ has eigenvalues $\{e^{-i\epsilon_{j}}\}$ where $\{\epsilon_{j}\}$ are no longer all real. Numerically, the $\epsilon_{j}$ with smallest (negative) imaginary part, which governs the slowest decaying eigenvector, was shown to describe $\bar{g}$ obtained from the purely unitary dynamics generated by $U_{F}$. More recently, it was argued in Ref. Mori, 2024 that in ergodic Floquet systems, if one adds a small amount of dephasing dissipation (see Fig. 1) and takes the system size to infinity first before taking the dissipation rate to zero, then the resulting dissipative Floquet operator has eigenvalues $e^{-i\epsilon_{j}}$ with an $\epsilon_{j}$ with smallest negative real part that matches well with the decay of autocorrelation functions (this order of limits has also been studied in the context of “anomalous relaxation”Sá et al. (2022); García-García et al. (2023); Yoshimura et al. (2023)). Ref. Mori, 2024 also gave some physical intuition for why $\bar{g}$ can be encoded as an eigenvalue of the dissipative Floquet operator: the unitary dynamics spreads an operator ballistically until it grows to a steady-state size $x^{*}\sim\frac{1}{\gamma}$, where dissipation and unitary evolution balance out. Then the dissipation decreases the amplitude of this operator by $e^{-\gamma x^{*}}\sim e^{-\bar{g}}$ where $\bar{g}\sim\mathcal{O}(1)$.
In this work, we use the aforementioned hydrodynamical picture for operator spreading to elucidate, unify, and generalize the approaches described above for obtaining $\bar{g}$. In particular, we explain in detail how the two coarse-graining prescriptions above affect operator growth, and directly relate the modified operator growth dynamics to autocorrelation functions from unitary evolution. For the weak dissipation approach, we derive the result $x^{*}\sim\frac{1}{\gamma}$ explicitly, and we show why the specific order of limits $\lim_{\gamma\rightarrow 0}\lim_{L\rightarrow\infty}$ is needed to extract $\overline{g}$. More generally, we show that in addition to the two coarse-graining procedures above, any coarse-graining procedure turning the unitary dynamics non-unitary by suppressing large operators is likely to encode $\bar{g}$ as an eigenvalue of the resulting Liouvillian.
Hydrodynamical description of operator spreading.—We will first review the hydrodynamical description of operator spreading in random unitary circuits derived in Refs. von Keyserlingk et al., 2018; Nahum et al., 2018. We will assume that the circuit has a brickwork geometry built out of two-site gates, acting on sites with local Hilbert space dimension $q$. Any operator can be written as a sum of $\mathbb{Z}_{q}$ Pauli strings, generated by $\mathbb{Z}_{q}$ clock and shift operators (from this point onward we will simply refer to these as operator strings). Therefore, we only need to consider the evolution of operator strings under the random unitary dynamics.
An operator string $\sigma^{\mu}$ generically evolves into a sum of operator strings: $\sigma^{\mu}(t)=\sum_{\nu}c_{\nu}^{\mu}(t)\sigma^{\nu}$ where $\nu$ runs over operator strings and $c_{\nu}^{\mu}(t)$ are complex coefficients. One important insight from Refs. von Keyserlingk et al., 2018; Nahum et al., 2018 is that the evolution of an operator under a layer of random unitaries roughly depends only on the left and right endpoints of the operator (the sites farthes to the left and right that the operator acts non-trivially). Therefore, we can group together all operator strings with the same left and right endpoints. For simplicity, we will let the original operator be at the left end of the system, so we roughly only need to consider the movement of the right endpoints of operators. [^2] Choosing the operator to start at the center of the system and/or keeping track of correlations between the left and right endpoints will only modify the results in this paper by subleading $\frac{1}{\mathrm{poly}(t)}$ factors. We denote the density of operators with right endpoint at site $x$ by
$\overline{\rho_{R}^{\mu}}(x,t)=\sum_{\nu}\overline{|c_{\nu}^{\mu}|^{2}}\delta(\mathrm{RHS}(\nu)=x)$ (2)
where we sum over operator strings $\sigma^{\nu}$ with right endpoint at $x$ [^3] We will use the same convention as in Ref. von Keyserlingk et al., 2018, where each unit cell (labeled by $x$) actually has two $q$ dimensional sites. It was shown in Ref. von Keyserlingk et al., 2018 that for random unitary circuits, $\overline{\rho_{R}^{\mu}}(x,t)$ takes the form of a propagating Gaussian with a peak moving at butterfly velocity $v_{B}$ that depends only on $q$ and variance growing linearly with $t$. This result was derived using the observation that $\overline{\rho_{R}}(x,t)$ satisfies the recurrence relation
$\begin{split}\overline{\rho_{R}}(x,t+1)&=2p(1-p)\overline{\rho_{R}}(x,t)+p^{2}\overline{\rho_{R}}(x-1,t)\\ &+(1-p)^{2}\overline{\rho_{R}}(x+1,t)\end{split}$ (3)
where $p=\frac{q^{2}}{q^{2}+1}$. This means that the right endpoints of operator strings perform a biased random walk, moving to the right with probability $p^{2}$, to the left with probability $(1-p)^{2}$, and staying on the same site otherwise. In this notation, $v_{B}=p^{2}-(1-p)^{2}=\frac{q^{2}-1}{q^{2}+1}$, which is smaller than the Lieb-Robinson velocity $1$ but approaches $1$ as $q\to\infty$.
We can package the above evolution of $\overline{\rho_{R}}(x,t)$ as follows. For every time step, we apply the matrix
$T(p)=\begin{pmatrix}1-p^{2}&(1-p)^{2}&0&\cdots\\ p^{2}&2p(1-p)&(1-p)^{2}&\cdots\\ 0&p^{2}&2p(1-p)&\cdots\\ 0&0&p^{2}&\cdots\\ \vdots&\vdots&\vdots&\vdots\end{pmatrix}$ (4)
onto the vector indexed by integer $x\in[1,L]$ with elements $\overline{\rho_{R}}(x,t)$. $T(p)$ is a stochastic matrix: the elements in every column sum to 1. This is because the unitarity of the evolution guarantees that $\mathrm{Tr}(\sigma^{\mu}(t)\sigma^{\mu}(t)^{\dagger})$ is conserved, so $\sum_{s}\rho_{R}^{\mu}(s)=1$ for all $t$.
The fact that $T(p)$ is a stochastic matrix means that it has an eigenvalue 1, with multiplicity 1. For Floquet systems and modified random circuits, the stochastic $L\times L$ matrix $T$ describing the evolution of $\overline{\rho_{R}}(x,t)$ may be different. We only require that $T$ is quasi-diagonal and that $T_{x,y}=T_{x+1,y+1}$ for $\mathcal{O}(1)<x,y<\mathcal{O}(L)$. The latter uniformity assumption corresponds to the statement that once an operator has grown large, but is still significantly smaller than the system size, how it grows subsequently does not depend on its current size.
Autocorrelation decay.$T(p)$ can be used to compute how autocorrelation functions decay. We will assume that under ergodic dynamics, $|c_{\nu}^{\mu}(t)|^{2}$ is the same for any two operators with the same left and right endpoints. In the infinite temperature state, [^4] We give a more detailed relation between infinite temperature correlation functions and product state correlation functions in the supplemental material sup we get
$\sqrt{|\langle\sigma^{\mu}(t)\sigma^{\mu}(0)\rangle|^{2}}=\sqrt{|c_{\mu}^{\mu}(t)|^{2}}$ (5)
We used the simple observation that the trace over any nontrivial operator string is zero. Note that $\overline{\rho}_{R}(n,t)\sim q^{2n}|c_{\mu}^{\mu}(t)|^{2}$ where $n$ is the extent of $\sigma^{\mu}$, so the autocorrelation function decays with the same rate as $\overline{\rho_{R}}(n,t)$. $\overline{\rho_{R}}(x,t)$ can be computed from $T(p)^{t}$ acting on a vector $|n\rangle$ whose only nonzero element is at site $n$. Therefore, to evaluate $\overline{\rho}_{n,t}$, we simply compute
$\overline{\rho}(n,t)=\langle n|T(p)^{t}|n\rangle=\sum_{k}\lambda_{k}^{t}\langle n|k_{R}\rangle\langle k_{L}|n\rangle$ (6)
where for the second equality, we decomposed $T(p)$ into left and right eigenvectors.
It may be tempting to immediately conclude that autocorrelations are governed by the largest eigenvalue $\lambda_{0}=1$ of $T(p)$. However, since $\langle n|k_{R}\rangle\langle k_{L}|n\rangle$ is exponentially small in $L$ for $k=0$ compared to $k\neq 0$, for times smaller than $\mathcal{O}(L)$, autocorrelation functions are actually governed by the subleading eigenvalues. We show in the supplemental material sup that there are $L-1$ eigenvalues $\lambda_{k}=2p(1-p)(1+\cos k)$ where $k=\frac{m\pi}{L}$ for $m=1,\cdots L-1$. These eigenvalues lead to a decay rate of $(4p(1-p))^{t}$ (up to $\frac{1}{\mathrm{poly}(t)}$ corrections) for $t<\mathcal{O}(L)$.
Finally at times $t\sim\mathcal{O}(L)$, the leading eigenvalue $\lambda_{0}=1$ dominates, causing autocorrelation functions to plateau. We can roughly estimate this plateau time by setting $4p(1-p)$ equal to the exponentially small value of $\langle n|0_{R}\rangle\langle 0_{L}|n\rangle$. For $n\sim\mathcal{O}(1)$, this givessup
$t_{\mathrm{plateau}}\sim-2L\frac{\log p(1-p)}{\log 4p(1-p)}$ (7)
$t_{\mathrm{plateau}}$ grows linearly with $L$, with the coefficient going to zero as $p\to 1$.
The above derivation of autocorrelation decay shows that there is no dependence on the observable $O$ or the initial state. Any local operator $O$ can be mapped onto another local operator by a finite number of layers of unitaries, so the long-time dynamics is unaffected by the choice of $O$. Furthermore, the computation of (1) in a product state is similar to the above computation; we simply need to also sum over $|c_{\mu}^{\nu}(t)|^{2}$ where $\sigma^{\nu}$ is a operator string for which the product state is an eigenstate. It turns out that this only changes the $\mathrm{poly}(t)$ factor sup ; the exponential decay rate remains the same.
More generally, as long as $T$ is stochastic, quasi-diagonal, and uniform, we expect that autocorrelation functions decay exponentially with a constant rate, up to times of order $L$. The decay rate given by subleading eigenvalues of $T$. In the above example, it is the second leading eigenvalue of $T$, but more generally there may be $\mathcal{O}(1)$ other eigenvalues (smaller than 1 but possibly larger than $e^{-\bar{g}}$) that also only appear only after $t\sim\mathcal{O}(L)$ because they are related to boundary condition of the biased walk at $x\sim L$. We give an example of this in the supplemental material sup .
We will now show that coarse-graining methods described earlier all have the effect of slightly modifying the matrix $T$. For the particular $T(p)$ in Eq. (4), they remove the eigenvalue 1 while, in appropriate limits, maintaining the eigenvalues around $4p(1-p)$. More generally, systems described by other choices of $T$, the coarse-graining methods remove the $\mathcal{O}(1)$ eigenvalues that appear at times $t\sim\mathcal{O}(L)$ so that the exponential decay observed for $t<\mathcal{O}(L)$ persists for all times. We will show that this implies that, for Floquet systems, the rate is encoded as an eigenvalue of the non-unitary generator of time translations.
Liouvillian gap: projection.— The superoperator $\mathbb{P}_{\ell}\mathbb{U}_{F}$ has eigenvalues within the unit circle, and it was conjectured in Refs. Prosen, 2002, 2004 that the leading such eigenvalue encodes $\bar{g}$. The projection $\mathbb{P}_{\ell}$ used to remove operators with size greater than $\ell$ can be easily implemented within the operator-hydrodynamical picture. We simply replace $T(p)$ by $P(\ell)T(p)$ where $P(\ell)$ is a $L\times L$ matrix $P(\ell)_{x,x}=1-\Theta(x,\ell)$ where $\Theta$ is the heaviside function. Clearly, since $P(\ell)T(p)$ is no longer stochastic, it does not have an eigenvalue $\lambda_{0}=1$. Notice that it is essentially a $T(p)$ on a smaller system size $\ell$, except with a modification at the $(\ell,\ell)$ matrix element that generically removes probability conservation. We show by solving the recurrence relation (2) with appropriate boundary conditions (see sup ) that $P(\ell)T(p)$ retains eigenvalues that approach $4p(1-p)$ in the thermodynamic limit. This means that autocorrelation functions decay as $(4p(1-p))^{t}$ even as $t\to\infty$; the projection removes the plateau at $t\sim\mathcal{O}(L)$. The eigenvectors corresponding to values near $4p(1-p)$ are all peaked near $x=\ell$. If an operator starts at size $1$, then $(P(\ell)T(p))^{t}$ drives it towards these dominant eigenvectors of $P(\ell)T(p)$ on an $O(\ell)$ timescale, and they dominate temporal correlations past that point.
While the hydrodynamic description usually does not apply after $t\sim\mathcal{O}(L)$ for Floquet systems (after this point, Floquet systems differ from RUCs, because they demonstrate a ramp and plateau in autocorrelation functions), the projection keeps the operators small compared to $L$ for all times. Since operators are kept small, the operator size distribution never reaches that required for the ramp and plateau. Operator hydrodynamics should apply to operators of size $\ell$ embedded in a much larger system of size $L$, so we can continue to use operator hydrodynamics at all times. Even in Floquet systems, we expect that the dominant eigenvectors of the dissipative evolution still correspond to operators of size near $\ell$, and a local operator will settle to them on an $O(\ell)$ timescale.
What does this mean for the spectrum of $\mathbb{P}_{\ell}\mathbb{U}_{F}$? Consider the dynamics of $(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}(O)$ for a local operator $O$. The hydrodynamical picture suggests that this operator will converge to the size $x\sim\ell$ eigenvectors over an $O(\ell)$ time-scale, and subsequently for all times decay as $\lambda^{t}$ for all times for some $\lambda$ (related to the aforementioned subleading eigenvalues of $T$). Therefore, for all times $t\geq O(\ell)$ we should have
$\left|\langle O\,(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}[O]\rangle\right|\sim\lambda^{t}$ (8)
Although the superoperator $\mathbb{P}_{\ell}\mathbb{U}_{F}$ has non-orthonormal left and right eigenvectors, at some $t$, the decay rate will still be dominated by the eigenvalue with the largest modulus. In other words, at some value of $t$, the eigenvalue modulus overcomes any transient features determined by wavefunction overlaps [^5] For example, after times $t\sim\mathcal{O}(1)$, the largest eigenvalue $\lambda_{0}=1$ of $T(p)$ begins to dominate. It follows that $\mathbb{P}_{\ell}\mathbb{U}_{F}$ needs to have a dominant eigenvalue $\lambda$. Now, notice that $\lambda$ implicitly depends on $\ell$; however we do not expect it to depend strongly on $\ell$. The idea is that autocorrelators for $\mathbb{U}_{F}^{t}$ can be represented as a sum over histories of Pauli matrices. $(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}$ involves the same histories, except those involving operators larger than $\ell$ are discarded. However arguments like those in Rakovszky et al. (2022); von Keyserlingk et al. (2022) suggest the contributions discarded are subleading, and suppressed in $\ell$. This suggests that $\langle O\,(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}[O]\rangle$ and $\langle O\,\mathbb{U}_{F}^{t}[O]\rangle=\langle O(t)O\rangle$ decay at similar exponential rates, so that $\lim_{\ell\rightarrow\infty}\lambda=e^{-\overline{g}}$. This establishes that the leading eigenvalue of $\mathbb{P}_{\ell}\mathbb{U}_{F}$ approaches $e^{-\overline{g}}$ as $\ell\rightarrow\infty$, assuming we take the thermodynamic $L\rightarrow\infty$ limit first.
Another way to see the independence of $\ell$ and the convergence to $e^{-\bar{g}}$ is the following. We do not expect $\left|\langle O\,(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}[O]\rangle\right|$ to depend strongly on $\ell$ for sufficiently large $\ell$ because the dominant eigenvectors of the evolution with $P_{\ell}$ are operators near size $\ell$, and according to our uniformity assumption, operators at size $\mathcal{O}(1)<\ell<\mathcal{O}(L)$ all behave in the same way under the hydrodynamical description. We expect this decay rate to converge to $e^{-\bar{g}}$ because we can choose $\ell$ to be very large, i.e. $\ell=\frac{1}{2}L$. Then $\mathbb{P}_{\ell}$ does not affect the dynamics at all up to $t\sim\mathcal{O}(L)$. According to our arguments regarding how $P(\ell)$ modifies the eigenvalue structure of $T$, this means that the same rate $\bar{g}$ describes how $\left|\langle O\,(\mathbb{P}_{\ell}\mathbb{U}_{F})^{t}[O]\rangle\right|$ decays both at times up to $\mathcal{O}(L)$ (where $\mathbb{P}_{\ell}$ does nothing) and at times much larger than $\mathcal{O}(L)$ (from which we extract $\lambda=e^{-\bar{g}}$).
Liouvillian gap: dissipation.— In the previous section we ‘coarse-grained’ the unitary evolution by projecting the dynamics onto the space of small operator strings; we refer to this procedure as hard truncation. In this section we show that the main conclusions from our previous section continue to hold when we more gently coarse-grain, by discarding longer operators through adding weak dissipation to the evolution. The effect of weak dissipation on the hydrodynamical description of operator spreading can also be implemented by a simple modification of $T(p)$. In the many body problem, interspersing unitary evolution with dissipation corresponds to stroboscopic evolution with $\mathbb{P}_{\gamma}\mathbb{U}_{F}$, with dissipative channel $\mathbb{P}_{\gamma}$. In the case where the dissipation is depolarising noise, for example, $\mathbb{P}_{\gamma}(\sigma^{\mu})=e^{-\gamma|\mu|}\sigma^{\mu}$, where $|\mu|$ is the Pauli-weight of $\mu$ and $\gamma$ is the strength of dissipation. In order to incorporate dissipative effects into operator hydrodynamics, we simply multiply $T(p)$ by a diagonal matrix $P(\gamma)$ which captures the effects of $\mathbb{P}_{\gamma}$ within the hydrodynamical picture. We take $P(\gamma)$ to be a diagonal matrix with $P(\gamma)_{x,x}=e^{-c\gamma x}$. Here, $c$ is a constant that depends on the particular channel. It does not change the result for $\gamma\to 0$, so for convenience in the following we will set $c=1$. $P(\gamma)$ captures the effects of dissipation because dissipation generically reduces operator amplitude, by an amount proportional to the operator weightSchuster and Yao (2023). [^6] For large $q$, the operator weight is approximately its extent, so we can roughly model dissipation at rate $\gamma$ by reducing the amplitude of operators by an amount exponential in their extent (the discrepancy between operator weight and operator extent can also be absorbed into the constant $c$, which does not affect the result as $\gamma\to 0$).
Similar to the hard truncation, the soft truncation described by $P(\gamma)$ also removes the eigenvalue 1 of $T(p)$. Furthermore, in the above limits, $[P(\gamma)T(p)]_{x,y}\sim T(p)_{x,y}$ for $x,y\ll\frac{1}{\gamma}$. We can use our results for $P(\ell)T(p)$ to show that the eigenvalues of $P(\gamma)T(p)$ are upper bounded by $4p(1-p)$. Specifically, $P(\ell)T(p)$ has the same eigenvalues as $P(\ell)\tilde{T}(p)$ where $\tilde{T}(p)$ is a symmetric real matrix (see sup ). Next, we can define $\tilde{T}_{\mathrm{pr}}(p)$ which has a maximum eigenvalue close to $\sim 4p(1-p)$ by $\tilde{T}(p)=\tilde{T}_{\mathrm{pr}}(p)-e^{-\gamma L}p^{2}|L\rangle\langle L|$. Since $P(\gamma),\tilde{T}(p)$, and $\tilde{T}_{\mathrm{pr}}(p)$ are all Hermitian and positive definite, and the spectrum of $P(\gamma)\tilde{T}(p)$ is the same as that of $P(\gamma)\tilde{T}_{\mathrm{pr}}(p)$ up to corrections $\sim\mathcal{O}(e^{-\gamma L})$, the spectrum of $P(\gamma)\tilde{T}(p)$ is bounded from above by the product of the largest eigenvalues of $P(\gamma)$ and $\tilde{T}_{\mathrm{pr}}(p)$, which is $e^{-\gamma}4p(1-p)$. We observe numerical evidence that the leading eigenvalue of $P(\gamma)T(p)$ indeed approaches $4p(1-p)$ as $L\to\infty,\gamma\to 0$ (see Fig. 2).
Furthermore, it is straightforward to see that the eigenvector $\overline{\rho_{R,k}}(x)$ of $P(\gamma)T(p)$ corresponding to eigenvalue $\lambda_{k}$ is peaked at size $-\frac{\log(\lambda_{k})}{\gamma}$, as argued by Ref. Mori, 2024. The eigenstate satisfies
$\begin{split}\lambda_{k}\overline{\rho_{R,k}}(x)&=e^{-\gamma x}2p(1-p)\overline{\rho_{R,k}}(x)+e^{-\gamma(x-1)}\overline{\rho_{R,k}}(x-1)\\ &+e^{-\gamma(x+1)}\overline{\rho_{R,k}}(x+1)\end{split}$ (9)
Suppose that $\overline{\rho_{R,k}}(x)$ is a smooth function of $x$ (we will show that this is true in an approximation of $P(\gamma)T(p)$ later in this section; there, $\overline{\rho_{R,k}}$ has a Gaussian-like shape). Then when $\overline{\rho_{R,k}}(x)$ peaks, it flattens out, so $\overline{\rho_{R,k}}(x^{*})\sim\overline{\rho_{R,k}}(x^{*}-1)\sim\overline{\rho_{R,k}}(x^{*}+1)$. If we in addition take $\gamma\to 0$, (9) gives $\lambda_{k}=e^{-\gamma x^{*}}$, so $x^{*}=-\frac{\log{\lambda_{k}}}{\gamma}$. Assuming that the leading eigenvalue of $P(\gamma)T(p)$ approaches $4p(1-p)$ in the limit $L\to\infty,\gamma\to 0$, we find that indeed operators get spread to a size $\sim-\frac{\log(4p(1-p))}{\gamma}$. Because the operator size distribution get frozen, and operators do not grow to size $\sim L$, our argument from the previous section applies to this case as well. Even for Floquet systems, where $T$ may be different from $T(p)$ in (4), we expect the same overall picture to hold: dissipation leads to operators growing ballistically to size $x_{*}=O(1/\gamma)$ and becoming stuck there subsequently, with their norm decaying $e^{-\gamma x_{*}t}$ ever after. It follows from a similar argument to that near Eq. (8) that the dissipative Floquet operator $\mathbb{P}_{\gamma}\mathbb{U}_{F}$ has a leading eigenvalue with modulus $e^{-\bar{g}}$.
To get intuition for the eigenvalues and eigenvectors of $P(\gamma)T(p)$, it is instructive to consider a simplified $T(p)$ obtained by dropping the upper diagonal terms. This reduces the second order recurrence relation to a first order recurrence relation, which makes it easily solvable. The resulting leading eigenvector and eigenvalues qualitatively behave in the same way as those of the full second order problem.
Expanding $T(p)$ for $p=1-\varepsilon$ in small $\varepsilon$ turns $T(p)$ into a lower triangular matrix $T(p)^{(1)}$. $P(\gamma)T(p)^{(1)}$ is still lower triangular, and its eigenvalues are given simply by its diagonal matrix elements. In addition to an eigenstate localized at $x=L$ with eigenvalue $\sim e^{-\gamma L}$, there is a second leading eigenvalue $\lambda_{1}=e^{-\gamma}2\varepsilon$. Other eigenvalues are smaller, at $e^{-\gamma x}2\varepsilon$. Therefore, if $\gamma L\ll 1$, the dominant eigenvalue corresponds to an eigenvector localized at $x=L$. However, if we take $L\to\infty$ first, and then $\gamma\to 0$, the leading eigenvalue of $P(\gamma)T(p)^{(1)}$ converges to $2\varepsilon$. For this simplified first order model, it is easy to see that autocorrelation functions also decay precisely as $(2\varepsilon)^{t}$. This is because if an operator grows, it cannot reduce back in size, and $2\varepsilon$ is the probability of remaining the same size.
To get the structure of the eigenstate, we can obtain all of the ratios $\frac{\overline{\rho_{R}}(x)}{\overline{\rho_{R}}(x-1)}$ (which is simple in this case because the recurrence relation is first order) and use a telescoping product to get
$\overline{\rho_{R}}=e^{\gamma x}\left(\frac{1-2\epsilon}{2\epsilon}\right)^{x}e^{-\frac{\gamma x(x+1)}{2}}\prod_{y<x}\frac{1}{1-e^{-\gamma(y-1)}}$ (10)
This state is roughly a Gaussian peaked at $x^{*}=-\frac{\log(2\epsilon)}{\gamma}$. It follows that $\gamma x^{*}=-\log(2\epsilon)=-\bar{g}$ for this simple model.
Figure 2: Leading eigenvalue of $P(\gamma)T(p)$ as a function of $\gamma$ for different system sizes, with $p=0.8$. Black dashed line is $4p(1-p)$.
Discussion.— Using an operator hydrodynamical picture, which is exact for RUCs and conjectured to work for ergodic Floquet circuits, we considered how operators grow in the presence of two very different forms of dissipation (hard versus soft truncation). We argued that at late times, a generic initial local operator tends towards a quasi-steady state with a profile peaked at size $x_{*}\sim\ell,1/\gamma$ in the hard, soft cases respectively (the latter result agrees with a prediction of Mori (2024)). In the RUC case, regardless of the choice of dissipation, the corresponding eigenvalue of the dissipative hydrodynamical process tends exactly to $e^{-\overline{g}}$ in the limit of weak dissipation, which controls the decay of autocorrelations. In the Floquet setting, we used the same operator spreading model to argue (less rigorously) that the leading eigenvalue of the dissipative many-body time translation generator $\mathbb{P}_{\ell}\mathbb{U}_{F}$ (or $\mathbb{P}_{\gamma}\mathbb{U}_{F}$) similarly determines the decay rate of autocorrelators in the weak dissipation limit.
We argued more generally that our results should be independent of the details of dissipation, provided it only suppresses large operators. This prediction comes from from the fact that our upper bound of the spectrum of $P(\gamma)T(p)$ does not depend on the particular form of $P(\gamma)$, as long as it is diagonal, is close to 1 for small $x$, and zero as $L\to\infty$.
There are some outstanding questions and comments. First of all, we have made the strong assumption that operator growth in ergodic Floquet many-body systems is governed by the same biased diffusive hydrodynamics as in the RUC case. This is certainly expected McCulloch and von Keyserlingk (2022); Zhou and Nahum (2020) but has not been proved rigorously. One might expect that if this conjecture does hold, then the leading eigenvalue of $P_{\ell}\mathbb{U}_{F}$, and hence $\overline{g}$, should simply be determined by the butterfly velocity ($v_{B}$), and front diffusion constant ($D$) of the corresponding hydrodynamical process. This is actually false; we give a family of models of biased diffusion in which have identical $v_{B},D$, but which predict different values of $\overline{g}$sup . Thus $\overline{g}$ is, if it can entirely be determined by the operator hydrodynamical equations, necessarily sensitive to details of the hydrodynamics which would usually be considered irrelevant.
On the practical side, it would be useful to find an efficient numerical method for determining the effective stochastic matrix $T$ given an ergodic unitary evolution, perhaps using some of the ideas in Refs. McCulloch and von Keyserlingk, 2022; Zhou and Nahum, 2020. There are other questions related to fundamental questions in the physics of ergodicity in quantum systems. For example, can one show that if all local operators decay exponentially, then the unitary dynamics has to have a biased random walk description like it does in RUCs? How are the results modified for ergodic Hamiltonian evolution, where the only conserved quantity is energy?
Finally, it would be interesting to see if similar ideas can be applied to classical chaos. Specifically, is there a useful analogue of operator hydrodynamics in classical dynamics? Observables on the classical phase space are functions, which can be decomposed into Fourier components. Similar to how a single operator string evolves into a sum of operator strings with growing size, a classical observable may grow under ergodic dynamics into a complicated sum with increasingly high frequency Fourier components. It would be interesting to check, for simple models of classical chaos, if this kind of observable growth in Fourier space has a simple biased random walk description like operator size in RUCs, and if properties of this random walk can be used to compute RP resonances. Furthermore, it has been shown that in single-particle quantum maps, a coarse-grained quantum Frobenius-Perron operator has eigenvalues that converge to the corresponding classical RP resonances in the semiclassical limit followed by the limit of weak coarse-graining. In contrast, for quantum mechanical models whose classical limits are not chaotic, all the eigenvalues go back to the unit circle in the said limits. It will be interesting to see if such dichotomy exists for many-body quantum systemsNonnenmacher (2003); García-Mata and Saraceno (2004).
Acknowledgements— We thank Sarang Gopalakrishnan, Tibor Rakovszky, Eric Schultz, Dries Sels, Sagar Vijay, and Zixia Wei for helpful conversations. C.Z. is supported by the Harvard Society of Fellows and the Simons Collaboration on Ultra Quantum Matter. CvK is supported by a UKRI Future Leaders Fellowship MR/T040947/2. This work was initiated and performed in part at the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452 and a grant from the Alfred P. Sloan Foundation (G-2024-22395).
Note added— We recently became aware of a parallel and independent work by Jacoby, Huse, and Gopalakrishnan Jacoby et al. which will appear on the same arXiv posting as this manuscript. Our conclusions appear to agree where they overlap.